### v2

Abstract
We discuss estimation and inference of conditional treatment effects in regression discontinuity designs with multiple scores. Aside from the commonly used local linear regression approach and a minimax-optimal estimator recently proposed by Imbens and Wager (2019), we consider two estimators based on random forests—honest regression forests and local linear forests—whose construction resembles that of standard local regressions, with theoretical validity following from results in Wager and Athey (2018) and Friedberg et al. (2020). We design a systematic Monte Carlo study with data generating processes built both from functional forms that we specify and from Wasserstein Generative Adversarial Networks that can closely mimic the observed data. We find that no single estimator dominates across all simulations: (i) local linear regressions perform well in univariate settings, but can undercover when multivariate scores are transformed into a univariate score—which is commonly done in practice—possibly due to the “zero-density” issue of the collapsed univariate score at the transformed cutoff; (ii) good performance of the minimax-optimal estimator depends on accurate estimation of a nuisance parameter and its current implementation only accepts up to two scores; (iii) forest-based estimators are not designed for estimation at boundary points and can suffer from bias in finite sample, but their flexibility in modeling multivariate scores opens the door to a wide range of empirical applications in multivariate regression discontinuity designs.
Keywords: Monte Carlo simulation; honest regression forests; local linear forests; nonparametric methods
1 Introduction
First introduced by Thistlethwaite and Campbell (1960) to study the effect of scholarship programs on subsequent academic outcomes, regression discontinuity (RD) designs are widely used to estimate causal effects in non-experimental settings. Sharp RD designs explore a particular treatment assignment mechanism in which individuals receive the treatment if an observed running variable scores above some known threshold. Under the potential outcomes framework, Hahn et al. (2001) show that the treatment is as good as randomly assigned in the population arbitrarily close to the assignment threshold if the conditional expectation functions (CEF) of the potential outcomes, with and without the treatment, are continuous at the cutoff. See, for example, Imbens and Lemieux (2008) and Cattaneo and Titiunik (2022) for early and recent reviews on RD designs.
Estimation and inference of treatment effects in this context are, therefore, a local problem in the neighborhood of the assignment cutoff. Following Hahn et al. (2001) and Porter (2003), local polynomial regressions remain the most popular estimation strategy in the empirical RD literature. At the heart of these local methods is bandwidth selection for the weighting kernel. In designs with a univariate running variable, Imbens and Kalyanaraman (2012) propose a data-driven algorithm that yields an optimal bandwidth by minimizing the asymptotic mean squared error (MSE) evaluated at the score cutoff. This procedure, however, leads to a non-negligible bias term that does not vanish in the limit. Calonico et al. (2014) put forth a bias-corrected estimator that first subtracts from the point estimate an estimated bias term and then adjusts the standard error accordingly so that valid inference can be conducted.
In many situations, the treatment status is determined by more than just one criterion. Yet univariate local linear regressions do not readily generalize to the case with multiple scores: kernel-based methods are less useful when the dimension exceeds two or three (see, for example, the discussion in Chapter 6 of Hastie et al., 2016). In addition, standard bias correction techniques for local linear regressions quickly become algebraically involved to derive and computationally difficult to implement in higher dimensions. Existing nonparametric estimators that work directly with multiple scores include methods proposed by Papay et al. (2011) and Choi and Lee (2018) that employ local linear regressions with multiple bandwidths selected by cross-validation. Zajonc (2012) extends the MSE-optimal bandwidth of Imbens and Kalyanaraman (2012) to the bivariate case. But, as pointed out by Imbens and Wager (2019), these methods yield oversmoothing bandwidths and asymptotically biased estimates, an issue also highlighted by Calonico et al. (2014). [^1] Choi and Lee (2018) also consider rule-of-thumb bandwidths that scale with $n^{-1/6}$, which are again oversmoothing in the sense that $nh^{5}\not\to 0$. To avoid the problem of selecting a multivariate bandwidth, Imbens and Wager (2019) propose an estimator based on numerical optimization that, given a second derivative bound $\mathcal{B}$ on the CEF of potential outcomes, is minimax-optimal among all estimators that are linear in the outcome variable over a class of CEF with second derivatives bounded by $\mathcal{B}$. This method, however, requires estimating the second derivative bound $\mathcal{B}$ and, as we show in Section 4, its performance depends crucially on accurate estimates of the nuisance parameter $\mathcal{B}$.
In this paper, we conduct a systematic Monte Carlo simulation to evaluate existing approaches to multivariate RD designs, including the minimax-optimal estimator of Imbens and Wager (2019) and the commonly used local linear regression approach with a transformed univariate score: to sidestep the complications of running local linear regressions in higher dimensions, empirical researchers often reduce the problem to a univariate one. For example, unemployment insurance programs usually set multiple eligibility criteria, such as age and prior labor force history in determining the maximum duration of benefits, but Schmieder et al. (2012) subset the data to consider only individuals with eligible employment histories within their respective age groups, and use age as the effective running variable. A similar subsetting approach has been taken by Jacob and Lefgren (2004) and Matsudaira (2008) to evaluate education policies where treatment eligibility is determined by standardized test scores on multiple subjects. Another common approach is to construct a univariate transformation of the multivariate score, especially in cases where the treatment boundary takes arbitrary shapes and finding a suitable subset of the data where a univariate score applies is infeasible. For example, to exploit geographic discontinuities, Black (1999) considers the shortest distance to the boundary of a geographic region; Keele and Titiunik (2015) transform bivariate coordinates to scalars by taking their rescaled Euclidean (chordal) distance to a given boundary point, and then run local linear regressions. Other examples include Battistin et al. (2009) and Clark and Martorell (2014) that use the minimum of multiple scores as the effective running variable in designs where the treatment status is determined by any of the scores crossing the threshold. These approaches, however, have several drawbacks. Considering only subsets of the data leaves unexplored a rich set of treatment effects that could be identified under a valid multivariate RD design, where the conditional average treatment effect as a multivariate function of any point along the treatment boundary is identified; see Section 2. Heterogeneous treatment effects of this kind can be important to answer policy questions. Besides, subsetting does not fully exploit signals from multiple dimensions and could suffer a loss of precision (Imbens and Wager, 2019).
Perhaps more precarious is the case where observations with the same distance to the cutoff of interest are collapsed to the same point on the real line: transforming a multivariate score may incur unexpected behaviors of the resulting density of the effective univariate score, violating key assumptions needed for both identification and estimation. In particular, the density of the transformed univariate score may be zero at the new cutoff. In Section 3.1.1, we formally discuss the issue of zero density and provide a characterization of bivariate densities with this problem. We also show in simulations that such problem can make local linear regressions biased and undercover the true treatment effect, even if the sample size gets large. Nevertheless, these aforementioned issues that could arise when reducing a multivariate RD design to a univariate one have not received much attention in empirical RD research and are difficult to detect in practice without knowing the true data generating process (DGP).
What should an empirical researcher do in the presence of multiple scores, then? We consider, as alternatives to these existing approaches, two machine-learning-type estimators that have gained popularity in estimating heterogeneous treatment effect under unconfoundedness, but have not been fully exploited in the empirical RD literature—honest regression forests of Wager and Athey (2018) and local linear forests of Friedberg et al. (2020). Like local linear regressions, these ensemble-tree methods construct a neighborhood near a given cutoff, weighting observations by how frequently they fall into the same leaf as the cutoff across all the trees in the forest. Regression forests return the weighted average of outcomes within that neighborhood as the estimate, whereas local linear forests use these forest weights to run a weighted linear regression to further capture any smooth signals. Thus, regression forests and local linear forests are, respectively, analogues of local constant and local linear regressions, with forest weights replacing kernel weights. Since the identified conditional treatment effect at a given cutoff—or more precisely, treatment boundary point, for multivariate designs—is simply the difference between two conditional means, we mimic the standard local linear regression estimation procedure by building two forests using treated and control observations, respectively, and obtain the estimated conditional treatment effect by subtracting the the control forest estimate from the treated forest estimate, both evaluated at the same treatment boundary point. These forest-based estimators are easy to implement, allow for valid inference, and can accommodate any fixed number of scores that does not grow with the sample size.
To evaluate these nonparametric estimators, we design a systematic Monte Carlo study with DGPs built both from functional forms that we specify, and, following the proposal of Athey et al. (2021), from Wasserstein Generative Adversarial Networks (WGANs) with a high degree of complexity that can closely mimic the observed data. We find that no single estimator dominates across all simulations: (i) local linear regressions perform well in univariate settings, but can undercover when multivariate scores are transformed into a univariate score, as is commonly done in practice, possibly due to the “zero-density” issue of the collapsed univariate score at the new cutoff; (ii) the minimax-optimal estimator has good bias-variance properties when the bound on the second derivative of the CEF is accurately estimated, but its performance is sensitive to the estimation of this nuisance parameter and its current implementation only accepts up to two scores; (iii) forest-based estimators are not designed for estimation at boundary points and can suffer from bias in finite sample—though local linear forests improve upon regression forests when the underlying signal exhibits strong linear trends—but their flexibility in modeling multivariate scores opens the door to a wide range of empirical applications in multivariate RD designs.
Our contribution is twofold. First, we point out a potential issue of the common practice of implementing univariate local linear regressions when the running variable is in fact multivariate, give a formal characterization of when such problem arises, and show in simulations that it can lead to biased estimates even for very large sample sizes. Second, we bring into play two additional valid, easy-to-implement estimators that can be leveraged in multivariate RD designs but have been largely overlooked in empirical RD research, and together with a systematic Monte Carlo evaluation of existing estimators, provide empirical researchers with a menu of nonparametric methods that they can tailor to different empirical settings.
The rest of the paper proceeds as follows. In Section 2, we begin with the identification of conditional treatment effects in multivariate RD designs. In Section 3, we review current estimation and inference practices using local linear regressions and the issue thereof, as well as the minimax-optimal estimator, followed by an introduction to honest regression forests and local linear forests. We present simulation results in Section 4. Section 5 concludes. Appendix A collects the proof of Proposition 1, and Appendix B provides additional information on the Monte Carlo study. All replication files can be accessed at github.com/yqi3/RDForest.
2 Setup and Identification
Suppose we have a sample of independent and identically distributed (i.i.d.) observations $\mathcal{S}_{n}=\{(Y_{i},X_{i})\}_{i=1}^{n}$, where $Y_{i}\in\mathbb{R}$ is the outcome of interest and $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d_{X}}$ is the $d_{X}$-dimensional running variable that determines the binary treatment status, $D_{i}\in\{0,1\}$, of the $i$-th unit according to some exogenous treatment assignment rule, $A:\mathcal{X}\to\{0,1\}$. If $d_{X}=1$, then $D_{i}=1$ if and only if $A(X_{i})=\mathds{1}\{X_{i}\geq x^{c}\}=1$, where $x^{c}$ is the cutoff as in the standard sharp RD design with a univariate score. In higher dimensions, the treatment assignment rule is more complex. For example, Keele and Titiunik (2015) study the effect of television advertising on election turnout, where the treatment status is determined by a geographic media market boundary. Following Zajonc (2012), let $\mathcal{T}:=\{x\in\mathcal{X}:A(x)=1\}$ denote the set of scores assigned to the treatment and let
:=$\mathscr{B}:=\{x\in\mathcal{X}:\forall\epsilon>0,\ \exists x_{0},x_{1}\in\mathrm{Ball}(x,\epsilon)\text{ such that }A(x_{0})=0\ \text{and}\ A(x_{1})=1\}$
denote the treatment boundary, i.e., a point $x\in\mathcal{X}$ lies on the treatment boundary if every $\epsilon$-ball around it contains both treated and control scores. In univariate designs, $\mathscr{B}$ is simply a singleton containing the one-dimensional cutoff.
Let $\bigl{(}Y_{i}(1),Y_{i}(0)\bigr{)}$ denote the pair of potential outcomes with and without the binary treatment. We are interested in the following conditional treatment effect at any $x^{c}\in\mathscr{B}$:
$\tau(x^{c}):=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|X_{i}=x^{c}].$
For each unit $i$, we only ever observe one of the potential outcomes. However, if we assume (i) the CEF $\mathbb{E}\left[Y_{i}(1)|X_{i}=x^{c}\right]$ and $\mathbb{E}\left[Y_{i}(0)|X_{i}=x^{c}\right]$ are continuous at all $x^{c}\in\mathscr{B}$, and (ii) $X$ has a density, $f_{X}(\cdot)$, that is bounded away from 0 in some neighborhood around $x^{c}$ for all $x^{c}\in\mathscr{B}$, then we can identify $\tau(x^{c})$ using neighboring observations around $x^{c}$:
For all $\text{For all $x^{c}\in\mathscr{B}$,}\quad\tau(x^{c})=$ $\lim_{x\to x^{c}:~{}A(x)=1}\mathbb{E}\left[Y_{i}(1)|X_{i}=x\right]-\lim_{x\to x^{c}:~{}A(x)=0}\mathbb{E}\left[Y_{i}(0)|X_{i}=x\right]$
$=$ $~{}\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=1\right]-\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=0\right],$ (1)
where $\tau(x^{c})$ is identified as the difference between the conditional expectation of the observed outcome for the treated population and the control population, respectively, both evaluated at $x^{c}$. [^2] A detailed proof, following from a generalization of the argument in Hahn et al. (2001), can be found in Zajonc (2012, p.56). Thus, under a valid multivariate RD design, we can identify the conditional treatment effect as a function of scores along the treatment boundary $\mathscr{B}$. Such heterogeneity may be of interest to the researcher. Alternatively, Zajonc (2012) and Keele and Titiunik (2015) discuss a summary measure of the average treatment effect along the boundary, i.e., $\mathbb{E}[Y_{i}(1)-Y_{i}(0)|X_{i}\in\mathscr{B}]$, which can be estimated given a sample of boundary points by averaging the estimated conditional treatment effects at those sample boundary points. However, as pointed out by Narita and Yata (2023), the Lebesgue measure of the boundary $\mathscr{B}$ is typically zero, and hence they consider the Hausdorff measure instead and propose a two-stage least squares estimator for the average treatment effect along the treatment boundary, weighted by the Hausdorff measure. In what follows, we focus on estimating conditional treatment effects at a given treatment boundary point $x^{c}\in\mathscr{B}$ and leave how exactly to aggregate these heterogeneous treatment effects to the empirical researcher.
3 Estimation and Inference
In this section, we review the current estimation and inference practices using local linear regressions and the minimax-optimal estimator of Imbens and Wager (2019) in multivariate RD designs. We highlight the problem of the common practice of transforming multivariate scores to a univariate one and then running local linear regressions. We conclude this section with an introduction to the honest regression forests of Wager and Athey (2018) and the local linear linear forests of Friedberg et al. (2020), and discuss how they can be leveraged to estimate heterogeneous treatment effects in RD designs with multiple scores.
3.1 Local Linear Regressions
Given a boundary point $x^{c}\in\mathscr{B}$, estimation of $\tau(x^{c})$ boils down to estimating two CEFs,
$\mu^{+}(x^{c})$ $:=\mathbb{E}\left[Y_{i}(1)|X_{i}=x^{c}\right]=\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=1\right],$
$\mu^{-}(x^{c})$ $:=\mathbb{E}\left[Y_{i}(0)|X_{i}=x^{c}\right]=\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=0\right].$
Consider univariate designs first. The standard estimation approach is to run a local linear regression on each side of the cutoff using some weighting kernel $K_{h}(\cdot)$ with bandwidth $h$,
$\widehat{\tau}_{\texttt{llr}}(x^{c};h):=$ $\widehat{\mu}_{\texttt{llr}}^{+}(x^{c};h)-\widehat{\mu}_{\texttt{llr}}^{-}(x^{c};h),$ (2)
$\widehat{\mu}_{\texttt{llr}}^{+}(x^{c};h):=$ $\bm{e}_{1}^{\intercal}\left(\arg\min_{(\beta_{0},\beta_{1})^{\intercal}}\sum_{i=1}^{n}\mathds{1}\{X_{i}\geq x^{c}\}K_{h}(X_{i}-x^{c})\left(Y_{i}-\beta_{0}-\beta_{1}(X_{i}-x^{c})\right)^{2}\right),$
$=$ $\bm{e}_{1}^{\intercal}\bigl{(}X_{+}^{\intercal}W(X_{+},x^{c};h)X_{+}\bigr{)}^{-1}X_{+}^{\intercal}W(X_{+},x^{c};h)Y_{+},$ (3)
$\widehat{\mu}_{\texttt{llr}}^{-}(x^{c};h):=$ $\bm{e}_{1}^{\intercal}\left(\arg\min_{(\beta_{0},\beta_{1})^{\intercal}}\sum_{i=1}^{n}\mathds{1}\{X_{i}<x^{c}\}K_{h}(X_{i}-x^{c})\left(Y_{i}-\beta_{0}-\beta_{1}(X_{i}-x^{c})\right)^{2}\right),$
$=$ $\bm{e}_{1}^{\intercal}\bigl{(}X_{-}^{\intercal}W(X_{-},x^{c};h)X_{-}\bigr{)}^{-1}X_{-}^{\intercal}W(X_{-},x^{c};h)Y_{-},$ (4)
[…]
Following Imbens and Kalyanaraman (2012), the choice of the bandwidth $h$ is usually obtained by minimizing the asymptotic expansion of
$\mathbb{E}\left[\bigl{(}\widehat{\tau}_{\texttt{llr}}(x^{c};h)-\tau(x^{c})\bigr{)}^{2}\right]$ (5)
with respect to $h$. However, Calonico et al. (2014) note that the MSE-optimal bandwidth obtained from optimizing (5), denoted as $h^{*}$, is oversmoothing in the sense that $n(h^{{*}})^{5}$ does not converge to 0, leading to a bias term that must be corrected in order to conduct valid inference. They propose subtracting from $\widehat{\tau}_{\texttt{llr}}(x^{c};h^{*})$ an estimated bias term and then adjust the standard error to account for the variation in the estimation of the bias term. This procedure, however, does not readily generalize to multivariate RD designs. Although Zajonc (2012) proposes a similar MSE-optimal bandwidth selector in the bivariate case, valid inference still requires bias correction due to oversmoothing. In empirical research involving multiple scores, the multivariate problem is usually reduced to one dimension so that univariate methods can apply.
3.1.1 Problem with Transforming Multivariate Scores
As mentioned earlier, common practices to reduce a multivariate RD design to a univariate one include subsetting the data and transforming the multivariate score to a one-dimensional running variable, and each has its own pitfalls. In particular, transforming a multivariate score could incur unexpected behaviors of the resulting density of the effective univariate running variable. In this section, we detail a situation that may arise under such transformation: the transformed univariate score may have zero density at the new cutoff, violating a key assumption for both the identification and estimation of treatment effects in RD designs.
In Section 4, we use real data from Keele and Titiunik (2015) to generate simulation data for the Monte Carlo study. Keele and Titiunik (2015) transform a bivariate score to a univariate one using rescaled Euclidean (chordal) distance to a given treatment boundary point and then estimate the treatment effect at the new cutoff $x^{c}=0$ using local linear regression with robust bias correction as in Calonico et al. (2014). The following proposition gives a general characterization of bivariate densities that have a zero density at the effective univariate cutoff $x^{c}=0$, once the bivariate score is collapsed to a univariate variable by taking the Euclidean distance between the bivariate score and a given boundary point.
Proposition 1. Let $(X_{1},X_{2})$ be a continuous bivariate random variable with joint density $f_{(X_{1},X_{2})}(x_{1},x_{2})$ over the support $\Omega(X_{1},X_{2})$. Given any point $(c_{1},c_{2})\in\Omega(X_{1},X_{2})$, consider the bivariate transformation $(X_{1},X_{2})\mapsto(E,V)$ where $E:=\sqrt{(X_{1}-c_{1})^{2}+(X_{2}-c_{2})^{2}}$ and $V:=X_{1}-c_{1}$ with joint support
$\Omega(E,V)=\left\{(e,v)\in\mathbb{R}^{2}:e\in\left[0,\max_{(x_{1},x_{2})\in\Omega(X_{1},X_{2})}\sqrt{(x_{1}-c_{1})^{2}+(x_{2}-c_{2})^{2}}\right],v\in[l(e),u(e)]\right\},$
where $l(e)$ and $u(e)$ are the lower and upper bound of $v$ that can depend on $e$ [^3] To simplify notation, we assume without loss of generality that the lower and upper bounds of $(e,v)$ are attained so that the joint support of $(E,V)$ is a closed subset of $\mathbb{R}^{2}$; otherwise the closed brackets need to be changed to open brackets accordingly. Then the Euclidean distance $E$ has a zero density at $e=0$, $f_{E}(0)=0$, if and only if
$\biggl{(}\left.\int_{l(e)}^{u(e)}\frac{e}{\sqrt{e^{2}-v^{2}}}\left\{f_{(X_{1},X_{2})}(v+c_{1},\sqrt{e^{2}-v^{2}}+c_{2})\right.\right.$
$\left.\left.\hskip 142.26378pt+f_{(X_{1},X_{2})}(v+c_{1},-\sqrt{e^{2}-v^{2}}+c_{2})\right\}dv\biggr{)}\right\rvert_{e=0}=0.$ (6)
We give the proof of Proposition 1 and two examples under which (6) holds in Appendix A. This characterization depends on the population joint density of the scores and is specific to the transformation chosen by the researcher. To illustrate, the first panel of Figure 1 is the true density of a transformed bivariate score that is uniformly distributed over $[-1,1]^{2}$, and the point to which we calculate the Euclidean distance is $(0,0)$; see Example A.1 for the derivation of this density. As in Keele and Titiunik (2015), we sign the Euclidean distance so that treated (control) observations receive a positive (negative) transformed score, and the new cutoff after transformation becomes 00, at which the density of the transformed score is exactly 00. In practice, the researcher has access to only finite samples, and problems of this kind are difficult to detect without knowing the true DGP. In the last two panels of Figure 1, we plot the raw histograms of the transformed bivariate score with respect to a given treatment boundary point using the exact same transformation employed by Keele and Titiunik (2015); each histogram uses data simulated from a different DGP built using the real data from Keele and Titiunik (2015), and thus the simulated bivariate scores do not follow any known distributions. [^4] We provide details on how the simulation data is generated in Section 4. In all three panels, the (empirical) density of the transformed bivariate score is zero at the effective univariate cutoff $x^{c}=0$. When such zero-density problem arises, not only is the conditional treatment effect at the transformed cutoff not identified because the conditional expectations are not defined, the problem is also reflected in estimation, especially for very large sample sizes, as we show in the simulations in Section 4.
Figure 1: True and empirical densities of transformed bivariate scores. From left to right: (i) True density of a bivariate uniform running variable on $[-1,1]^{2}$ transformed using signed Euclidean distance; (ii) histogram of transformed bivariate coordinates using signed chordal distance to the selected boundary point, plotted using all $40$ million observations drawn from the WGAN-generated DGP; (iii) histogram of transformed bivariate coordinates using signed chordal distance to the selected boundary point, plotted using a sample of $1$ million observations from the polynomial DGP. The number of bins for both histograms is 10,000.
3.2 A Minimax-Optimal Estimator
To avoid the problem of choosing a multivariate bandwidth, Imbens and Wager (2019) propose an estimator based on numerical optimization that is minimax-optimal among all estimators that are linear in the outcome variable over all problems with $\sigma_{i}^{2}:=Var[Y_{i}|X_{i}]$ and a second derivative bound $\mathcal{B}$ on the CEF of potential outcomes. Specifically, their estimator for $\tau(x^{c})$ takes the form
$\widehat{\tau}_{\texttt{mm}}(x^{c};\mathcal{B}):=\sum_{i=1}^{n}\widehat{\gamma}_{i}(x^{c};\mathcal{B})Y_{i},$ (7)
where
$\widehat{\gamma}(x^{c};\mathcal{B})$ $=\arg\min_{\gamma}\left\{\sum_{i=1}^{n}\gamma_{i}^{2}\sigma_{i}^{2}+b^{2}(\gamma;\mathcal{B})\right\},$ (8)
$b(\gamma;\mathcal{B})$ $:=\sup_{\mu^{+},\mu^{-}}\left\{\sum_{i=1}^{n}\gamma_{i}\mu_{A_{i}}(X_{i})-(\mu^{+}(x^{c})-\mu^{-}(x^{c})):\forall\,x,\lVert\nabla^{2}\mu^{+}(x)\rVert\leq\mathcal{B},\lVert\nabla^{2}\mu^{-}(x)\rVert\leq\mathcal{B}\right\},$
[…]
$\mathcal{I}_{\alpha}:\liminf_{n\to\infty}\inf\left\{\mathbb{P}[\tau(x^{c})\in\mathcal{I}_{\alpha}]:\forall\,x,\lVert\nabla^{2}\mu^{+}(x)\rVert\leq\mathcal{B}\text{ and }\lVert\nabla^{2}\mu^{-}(x)\rVert\leq\mathcal{B}\right\}\geq 1-\alpha.$ (9)
Implementing the minimax-optimal estimator (7) requires estimates of $\sigma_{i}^{2}$ and $\mathcal{B}$. Imbens and Wager (2019) recommend estimating $\sigma_{i}^{2}$ by averaging the residuals from an ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $A(X_{i})$. Estimating $\mathcal{B}$, however, is problem-specific and less straightforward. Imbens and Wager (2019, p.273) recommend estimating the CEF globally using, for example, second-order polynomials, and then multiplying the maximal estimated curvature by 2 or 4. In one of our simulations in Section 4 using data from Lee (2008), we find that using second-order polynomials yields estimates of $\mathcal{B}$ that are too small, and multiplying by $4$ does not help; only after multiplying the maximal estimated curvature by $30$ do we get an estimate that is close to the true (least upper) bound $\mathcal{B}$.
3.3 Honest Regression Forests and Local Linear Forests
We now introduce two additional theoretically valid estimators that can be leveraged in multivariate RD designs but have been largely overlooked in the literature—the honest regression forests of Wager and Athey (2018) and the local linear forests of Friedberg et al. (2020). Both estimators are variants of random forests originally proposed by Breiman (2001) and have received popularity in estimating heterogeneous treatment effects under unconfoundedness.
3.3.1 Honest Regression Forests
A regression forest evaluated at some point $x\in\mathcal{X}$, denoted as $RF(x)$, is an average of $B$ trees, $\{T_{b}(x)\}_{b=1}^{B}$, each fitted using a bootstrapped sample or a random subsample of $\mathcal{S}_{n}:=\{(Y_{i},X_{i})\}_{i=1}^{n}$. Following Athey et al. (2019), Friedberg et al. (2020), and Wager and Athey (2018), we focus on forests built using random subsamples, for which central limit theorems are established. Given a subsample of size $s$ from $\mathcal{S}_{n}$, $\{(Y_{i_{k}},X_{i_{k}})\}_{k=1}^{s}$, a regression tree recursively divides the sample into mutually exclusive and collectively exhaustive convex subsets called leaves. For a point $x\in\mathcal{X}$, let $L(x)$ denote the leaf that contains $x$. A tree $T$ estimates the CEF, $\mathbb{E}[Y_{i}|X_{i}=x]$, by averaging the observations that fall into the same leaf as $x$:
$T(x)=\frac{1}{|L(x)|}\sum_{i_{k}:X_{i_{k}}\in L(x)}Y_{i_{k}}.$
Similar to selecting an MSE-optimal bandwidth for local linear regressions, we look for partitions of the feature space that are MSE-optimal. Unlike local linear regressions, trees do not target the MSE at a particular treatment boundary point, nor do they have an analytical form for the asymptotic expansion of the overall MSE,
$\mathbb{E}\left[\bigl{(}Y_{i}-T(X_{i})\bigr{)}^{2}\right],$ (10)
and searching for all possible partitions is computationally infeasible. Instead, standard splitting rules rely on greedy algorithms that immediately improve the in-sample fit of leaf-wide averages after each split.
Estimates from a single tree can be highly unstable without regularization. In the extreme case, one can achieve perfect in-sample fit by partitioning the sample data so that each leaf contains only one observation. Regression forests stabilize predictions from individual trees by averaging many of them: for a collection of $b=1,...,B$ trees, we fit each tree $T_{b}(x)$ using a random subsample of size $s$, and obtain the regression forest estimate by averaging across all $B$ trees, $\frac{1}{B}\sum_{b=1}^{B}T_{b}(x)$. To achieve a sizable reduction in variance, we may want to decorrelate the trees as much as possible by, for example, randomly selecting a dimension to split at each node. We assume throughout that $B$ is large enough so that the Monte Carlo variability of which subsamples are drawn is negligible and the following definition of regression forests from Wager and Athey (2018) given the training sample $\mathcal{S}_{n}$ and subsample size $s$ applies:
$RF\bigl{(}x;s,\mathcal{S}_{n}\bigr{)}:={\binom{n}{s}}^{-1}\sum_{1\leq i_{1}<...<i_{s}\leq n}\mathbb{E}_{\xi}\bigl{[}T(x;\xi,\{(Y_{i_{k}},X_{i_{k}})\}_{k=1}^{s})\bigr{]},$ (11)
which averages across trees fitted using all possible draws of a size-$s$ subsample, marginalized over an auxiliary random term $\xi$ that serves as a decorrelation tool encoding, for example, the random split information. To establish asymptotic normality of regression forests, Wager and Athey (2018) leverage a specific subsampling scheme satisfying a property called honesty to reduce bias, which requires that the subsample used to make the splits is independent of the subsample used to estimate the within-leaf conditional means. We refer the reader to Wager and Athey (2018) for a discussion on honesty and other regularity conditions needed to establish asymptotic normality of honest regression forests.
We now define the estimator based on honest regression forests for the conditional treatment effect at some point $x^{c}\in\mathscr{B}$ along the treatment boundary, $\tau(x^{c})$:
$\widehat{\tau}_{\texttt{rf}}(x^{c})$ $:=\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})-\widehat{\mu}_{\texttt{rf}}^{-}(x^{c}),$ (12)
$\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})$ $:=RF\bigl{(}x^{c};s_{1},\{(Y_{i},X_{i})\in\mathcal{S}_{n}:A(X_{i})=1\}\bigr{)},$
$\widehat{\mu}_{\texttt{rf}}^{-}(x^{c})$ $:=RF\bigl{(}x^{c};s_{0},\{(Y_{i},X_{i})\in\mathcal{S}_{n}:A(X_{i})=0\}\bigr{)}.$
Just like estimation using local linear regressions, $\widehat{\tau}_{\texttt{rf}}(x^{c})$ takes the difference between a regression forest fitted using observations that fall into the treated region and a forest fitted using observations that fall into the control region, both evaluated at the same treatment boundary point $x^{c}\in\mathscr{B}$.
Remark 1. Although it seems that the local linear estimator (2) depends on one more parameter, $h$, than the regression forest estimator (12), we note that this is not the case. Each tree $T$ in the regression forest implicitly defines a neighborhood $L(x^{c})$ for the treatment boundary point $x^{c}$. These neighborhoods are then aggregated, producing a weight for each observation that is proportional to the fraction of time it falls into the same leaf as $x^{c}$ across all trees in the regression forest. These forest weights are analogous to the weights generated by a kernel with bandwidth $h$. Thus (12) implicitly depends on some “pseudo-bandwidth” generated by tree partitions across the forest. Analogously, the $\gamma$ weights of the minimax-optimal estimator (7) play the role of kernels and bandwidths.
We can equivalently express regression forests as solutions to weighted least-squares constant regressions:
$\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})$ $=\arg\min_{\theta}\sum_{i=1}^{n}D_{i}W_{\texttt{rf}}^{+}(X_{i},x^{c})\left(Y_{i}-\theta\right)^{2},$
$=\left(\mathbf{1}_{n_{1}}^{\intercal}W_{\texttt{rf}}^{+}(X_{+},x^{c})\mathbf{1}_{n_{1}}\right)^{-1}\mathbf{1}_{n_{1}}^{\intercal}W_{\texttt{rf}}^{+}(X_{+},x^{c})Y_{+},$ (13)
$\widehat{\mu}_{\texttt{rf}}^{-}(x^{c})$ $=\arg\min_{\theta}\sum_{i=1}^{n}(1-D_{i})W_{\texttt{rf}}^{-}(X_{i},x^{c})\left(Y_{i}-\theta\right)^{2},$
$=\left(\mathbf{1}_{n_{0}}^{\intercal}W_{\texttt{rf}}^{-}(X_{-},x^{c})\mathbf{1}_{n_{0}}\right)^{-1}\mathbf{1}_{n_{0}}^{\intercal}W_{\texttt{rf}}^{-}(X_{-},x^{c})Y_{-},$ (14)
where the elements in (13) and (14) are analogous to the terms in (3) and (4), except $X_{+}$ and $X_{-}$ are now $n_{1}\times 1$ and $n_{0}\times 1$ vectors of ones, denoted respectively by $\mathbf{1}_{n_{1}}$ and $\mathbf{1}_{n_{0}}$, and we replace the kernel weights with forest weights,
$W_{\texttt{rf}}^{+}(X_{i},x^{c}):=\frac{1}{B}\sum_{b=1}^{B}\frac{\mathds{1}\{X_{i}\in L^{+}_{b}(x^{c})\}}{|L^{+}_{b}(x^{c})|},$ (15)
where $L^{+}_{b}(x^{c})$ denotes the leaf containing $x^{c}$ in the $b$-th tree fitted using the treated observations; $W_{\texttt{rf}}^{-}$ is defined similarly for the control observations. More generally, the least-square representation of forests in (13) and (14) effectively views $\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})$ and $\widehat{\mu}_{\texttt{rf}}^{-}(x^{c})$ as solutions to the local moment conditions,
$\mathbb{E}\left[(Y_{i}-\theta)^{2}|X_{i}=x^{c},D_{i}=1\right]=0\quad\text{and}\quad\mathbb{E}\left[(Y_{i}-\theta)^{2}|X_{i}=x^{c},D_{i}=0\right]=0,$
and solves for $\theta$ using their empirical analogues with forest weights that reflect the importance of each observation $i$ in terms of minimizing the objective function. Athey et al. (2019) use forests as weighting kernels to solve parameters identified by local moments in a more general setting.
3.3.2 Honest Local Linear Forests
Though a powerful nonparametric tool and popular in applications, regression forests resemble local constant regressions with convex weights, and hence cannot extrapolate and are known to have limited ability to capture smoothness or linear trends in the CEF, as noted by Friedberg et al. (2020) and the references therein, especially at points near the boundary of the support where we observe data asymmetrically from only one side of a given boundary point. To improve the performance of forests in the presence of strong linear signals, Friedberg et al. (2020) build on the honest regression forests of Wager and Athey (2018) and propose a linear regression adjustment by running a ridge local linear regression with forest weights, under which we have the following estimator for $\tau(x^{c})$:
$\widehat{\tau}_{\texttt{llf}}(x^{c})$ $:=\widehat{\mu}_{\texttt{llf}}^{+}(x^{c})-\widehat{\mu}_{\texttt{llf}}^{-}(x^{c}),$
$\widehat{\mu}_{\texttt{llf}}^{+}(x^{c})$ $:=\bm{e}_{1}^{\intercal}\left(\arg\min_{(\beta_{0},\beta_{1})^{\intercal}}\sum_{i=1}^{n}D_{i}W_{\texttt{rf}}^{+}(X_{i},x^{c})\bigl{(}Y_{i}-\beta_{0}-\beta_{1}^{\intercal}(X_{i}-x^{c})\bigr{)}^{2}+\lambda\|\beta_{1}\|_{2}^{2}\right),$ (16)
$\widehat{\mu}_{\texttt{llf}}^{-}(x^{c})$ $:=\bm{e}_{1}^{\intercal}\left(\arg\min_{(\beta_{0},\beta_{1})^{\intercal}}\sum_{i=1}^{n}(1-D_{i})W_{\texttt{rf}}^{-}(X_{i},x^{c})\bigl{(}Y_{i}-\beta_{0}-\beta_{1}^{\intercal}(X_{i}-x^{c})\bigr{)}^{2}+\lambda\|\beta_{1}\|_{2}^{2}\right),$ (17)
where, similar to the expressions in (3) and (4) for local linear regressions, $\bm{e}_{1}$ is the standard basis in $\mathbb{R}^{(d_{X}+1)}$ with the first element being $1$ and 00 elsewhere to pick out the fitted intercept $\beta_{0}$; $W_{\texttt{rf}}^{+}(X_{i},x^{c})$ and $W_{\texttt{rf}}^{-}(X_{i},x^{c})$ are the forest weights defined in (15), [^5] To account for the fact that local linear forests will use linear regression adjustments to estimate conditional means (as opposed to using simple weighted means, as in the case of regression forests), Friedberg et al. (2020) also modify the standard splitting rules described in Section 3.3.1 to “ridge residual splitting;” see Section 2.1 of Friedberg et al. (2020). effectively replacing the kernel weights in (3) and (4); $\lambda\|\beta_{1}\|_{2}^{2}$ is the ridge penalty term that prevents overfitting to local trends, and as explained in Friedberg et al. (2020), is needed both to improve finite-sample performance and for proving convergence rates to establish a central limit theorem for local linear forests. Just like local linear regressions and regression forests, local linear forests (16) and (17) can also be expressed as solutions to weighted least-square problems; see equation (5) of Friedberg et al. (2020, p. 507).
The assumptions needed for the asymptotic normality of local linear forests overlap to a large extent with the conditions required for honest regression forests to be asymptotically normal. However, with an additional assumption on the smoothness of the CEF $\mathbb{E}[Y_{i}|X_{i}=x]$—that it is differentiable with a Lipschitz continuous derivative, as opposed to the Wager and Athey (2018) condition that only requires it to be Lipschitz continuous—Friedberg et al. (2020) show an improved convergence rate of local linear forests relative to the rate of regression forests, effectively exploiting the additional smoothness assumption with a linear adjustment. We refer the reader to Wager and Athey (2018) and Friedberg et al. (2020) for details on the technical assumptions and discussions thereof. Pointwise confidence intervals for both honest regression forests and honest local linear forests are constructed using variance estimates from the bootstrap of little bags algorithm of Sexton and Laake (2009), with consistency results established by the random forest delta method of Athey et al. (2019, Theorem 6). In our Monte Carlo study, we find that local linear forests are in general less biased than regression forests and have better coverage rates whenever the simulation DGP exhibits strong linear trends.
Remark 2. One caveat of using local linear forests in multivariate RD designs is that the point at which we evaluate local linear forests, $x^{c}\in\mathscr{B}$, is necessarily a boundary point of the support of $X_{i}$ for both treated and control populations. On the other hand, the central limit theorem of Friedberg et al. (2020, Theorem 1) only applies to points in the interior of the support, as the proof starts with a usual Taylor expansion of $\mathbb{E}[Y_{i}|X_{i}=x]$ around $x^{c}$. [^6] Asymptotic normality of local linear regressions also relies on a similar Taylor expansion, but it is the CEF of potential outcomes that gets expanded (as opposed to the CEF of observed outcomes conditional on either the treated or the control population), and so in this case $x^{c}\in\mathscr{B}$ is an interior point of the support of $X$. We sidestep this problem by adding an arbitrarily small “buffer” to $x^{c}\in\mathscr{B}$: we pick another point $\tilde{x}^{c}$ in the interior of the treated score region such that $\|\tilde{x}^{c}-x^{c}\|=\epsilon$ for an arbitrarily small $\epsilon>0$ and evaluate the treated local linear forest $\widehat{\mu}^{+}_{\texttt{llf}}$ at $\tilde{x}^{c}$ to approximate $\widehat{\mu}^{+}_{\texttt{llf}}(x^{c})$; we do the same for the control local linear forest $\widehat{\mu}^{-}_{\texttt{llf}}$. We require $\mu^{+}(x^{c})$ and $\mu^{-}(x^{c})$ to satisfy the assumptions in Friedberg et al. (2020, Theorem 1), which imply that both $\mu^{+}(x^{c})$ and $\mu^{-}(x^{c})$ are Lipschitz in $x^{c}$ and therefore very close to $\mu^{+}(\tilde{x}^{c})$ and $\mu^{-}(\tilde{x}^{c})$, respectively, for small $\epsilon$. Ideally, one should adjust the confidence interval accordingly to account for such approximation, but in our simulations we choose $\epsilon=10^{-30}$ so that such correction is negligible up to machine precision.
4 Simulations
In this section, we conduct a systematic Monte Carlo study to evaluate the two existing estimation and inference practices in multivariate RD designs, as well as the two forest-based estimators that have not received much attention in the empirical RD literature before. All replication files can be accessed at github.com/yqi3/RDForest.
4.1 Data
We first consider a univariate design simulated using data from Lee (2008), as the Lee (2008) data is a popular choice to conduct simulations in the RD literature. Lee (2008) studies the effect of being the current incumbent party in a district on the votes obtained in the district’s next election. The running variable is univariate and measures the difference in vote share between the Democratic candidate and their strongest opponent in the current election, and the outcome variable is the Democratic vote share in the next election. This univariate design normalizes the assignment cutoff to 00, and units with $X_{i}\geq 0$ are treated (i.e., winning the current election).
We then consider a bivariate design simulated using data from Keele and Titiunik (2015), who study the effect of television advertising on election turnout, where the running variable is the two-dimensional geographic coordinate of a voter’s location and the outcome of interest is election turnout and takes binary values. Treatment status is determined by a geographic media market boundary. Their data also contains several other outcome variables for placebo analysis. In addition to election turnout, we include age and housing price per square foot in our simulation to cover both discrete and continuous outcome variables. Keele and Titiunik (2015) consider three equally spaced points along the treatment boundary, depicted as black squares in Figure 3, to be the points at which they estimate the treatment effect Without loss of generality, we pick the middle point, circled in green in Figure 3, to be the boundary point of interest in our simulation.
We consider designs up to two scores in our Monte Carlo study because the current R implementation of the minimax-optimal estimator $\widehat{\tau}_{\texttt{mm}}(x^{c};\mathcal{B})$ in (7) provided by Imbens and Wager (2019) only accepts up to two scores. But in work in progress, we illustrate how to use the two forest-based estimators in RD designs with more than two scores to explore heterogeneous treatment effects along the treatment boundary in an empirical application.
4.2 Generating Simulation Data
To evaluate the aforementioned nonparametric estimators, we design a systematic Monte Carlo study with data generating processes built both from functional forms that we specify, and, following the proposal of Athey et al. (2021), from Wasserstein Generative Adversarial Networks (WGANs) with a high degree of complexity that can closely mimic the observed data.
We first consider fitting the observed data with a polynomial, which is a popular choice in the RD literature to evaluate estimators. Imbens and Kalyanaraman (2012), Calonico et al. (2014), and Calonico et al. (2019) all consider simulation data generated by using the Lee (2008) data to fit two $5$-th order global polynomials separately for the treated and control populations, which we also consider in our simulation. For the Keele and Titiunik (2015) data, we fit a $3$-rd order polynomial for continuous outcomes (i.e., age and housing price) and a logistic regression using, again, a $3$-rd order polynomial inside the logit link function for the binary outcome turnout, separately for the treated and control populations. [^7] We also tried higher order polynomials and the results do not qualitatively differ. Details on these polynomial DGPs can be found in Appendix B.1.
As argued by Athey et al. (2021), Monte Carlo studies are more credible when the data used to evaluate the performance of econometric methods does not come from researcher-chosen functional forms. Following their proposal, we also consider simulation data generated from WGAN, which trains two neural networks—one called the generator and the other called the critic—in an adversarial setting to minimize the Wasserstein distance between a real dataset and the artificial dataset generated from the neural nets. We refer the reader to Athey et al. (2021) for a technical introduction to WGAN.
Using WGAN to generate data to evaluate estimators for treatment effects, however, is problem-specific, since we never directly observe the treatment effect in the data, but rather we “identify” it under a set of identification assumptions. Athey et al. (2021) provide an example of how to generate simulation data to evaluate estimators of treatment effects under unconfoundedness, where they train two conditional WGANs to learn $X_{i}|D_{i}$ and $Y_{i}|(X_{i},D_{i})$ from the real sample. They then draw a fake sample of treatment status and covariates from the WGAN-generated distribution of $X_{i}|D_{i}$, and for each $(X_{i},D_{i})$ drawn, an outcome is drawn from the WGAN-generated distribution of $Y_{i}|(X_{i},D_{i})$. Under unconfoundedness, $Y_{i}|(X_{i},D_{i}=d)$ identifies $Y_{i}(d)|X_{i}$ and hence for each $Y_{i}$ drawn, its counterfactual outcome is generated by mutating the corresponding treatment status $D_{i}$ to $(1-D_{i})$ but keeping the same $X_{i}$, i.e., the counterfactual outcome is drawn from the WGAN-generated distribution of $Y_{i}|(X_{i},1-D_{i})$. The true average treatment effect is simulated by averaging the difference between the pair of generated outcome and its counterfactual over a large WGAN-generated sample.
Our setting, with a different design and identification assumption for the treatment effect of interest, requires a different way of using WGAN to generate suitable simulation data. Specifically, the treatment effects are only identified for $X_{i}\in\mathscr{B}$, so we need to simulate the treatment effect at each treatment boundary point of interest, instead of simply taking averages as in the unconfoundedness setting. To do so, we first train two conditional WGANs to learn $X_{i}|D_{i}$ and $Y_{i}|(X_{i},D_{i})$ from the real sample, as in Athey et al. (2021). But for a given treatment boundary point $x^{c}\in\mathscr{B}$, we draw a separate artificial sample to simulate the true treatment effect at $x^{c}$: after training the WGAN to learn the distribution of $Y_{i}|(X_{i},D_{i})$, we generate 100,000 outcomes from $Y_{i}|(X_{i}=x^{c},D_{i}=1)$ and another 100,000 outcomes from $Y_{i}|(X_{i}=x^{c},D_{i}=0)$, and take the average of their differences as the true treatment effect at $x^{c}$. [^8] We note that WGANs do not directly yield closed-form estimates of the fitted distribution, but rather generate a new, artificial sample of arbitrary size that closely mimics the real sample. This is why we need to simulate the true treatment effect at a given treatment boundary point $x^{c}\in\mathscr{B}$, unlike in the polynomial case where we can directly calculate the true treatment effects from the coefficients of the fitted polynomials evaluated at $x^{c}$. Figures 2 and 3 visualize the actual data and the simulated data from the polynomial specification and WGAN. In the Appendix, we include additional visualization of the WGAN-generated DGPs in Figures B.1B.7, and we report in Table B.2 the Wasserstein distance between the real data and the WGAN data.
Figure 2: Lee (2008): Democratic victory margin in the current election against the vote share in the next election. Panel A shows the observed data. Panel B plots the model data fitted using a 5-th order polynomial with different coefficients for the treated and control observations. Panel C plots the WGAN data. The cutoff is normalized to 0. The figure is generated using the rdrobust package (Calonico et al. (2014)) with default smoothing.
We use the default options available from the wgan package of Athey et al. (2021) to train the WGANs; see Appendix B.3 for details. Following Section 5 of Athey et al. (2021), we conduct a thorough robustness check to assess the variability of the Monte Carlo results with respect to sampling variation and WGAN tuning parameters, with robustness results reported in Appendix B.5.
Figure 3: Visualization of the sample data in Keele and Titiunik (2015) (left) as well as simulated polynomial and WGAN DGPs (middle and right, respectively). The outcome variable is the binary election turnout. Coordinates below the black curve (treatment boundary) are treated, and the green circle shows the boundary point at which we estimate the treatment effect. To visualize the polynomial DGPs, 2,500 coordinates are sampled with replacement from the original sample plus a noise term $N(0,0.01^{2})$.
4.3 Monte Carlo Results
We report the Monte Carlo simulation results in Table 1. All simulations are implemented in R, where we use the rdrobust package of Calonico et al. (2014) for local linear regressions, the optrdd package of Imbens and Wager (2019) for the minimax-optimal estimator, and the grf package of Athey et al. (2019) for honest regression forests and local linear forests—all with default options without tuning, except we specify a few hyperparameters for the two forest-based estimators for reasons we explain in Appendix B.6.
Local Linear Regression Minimax-Optimal Honest Regression Forest Honest Local Linear Forest
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) (14) (15) (16) (17) (18) (19) (20)
Sample size $n=$ $1,000$ $5,000$ $10,000$ $50,000$ $100,000$ $1,000$ $5,000$ $10,000$ $50,000$ $100,000$ $1,000$ $5,000$ $10,000$ $50,000$ $100,000$ $1,000$ $5,000$ $10,000$ $50,000$ $100,000$
Univariate Design: Polynomial DGP
Panel A: Vote Share (truth $\bm{=0.0400}$)
Bias 0.0161 0.0116 0.0078 0.0026 0.0011 0.0378 0.0257 0.0211 0.0134 0.0102 0.0344 0.0072 0.0043 0.0006 0.0019 0.0180 0.0051 0.0036 0.0008 0.0019
Variance 0.0021 0.0005 0.0003 0.0001 0.0000 0.0008 0.0002 0.0001 0.0000 0.0000 0.0014 0.0013 0.0013 0.0015 0.0013 0.0030 0.0020 0.0019 0.0017 0.0014
95% coverage 91.1% 86.8% 89.1% 91.2% 92.9% 80.5% 68.7% 64.8% 49.8% 51.6% 85.1% 94.9% 95.6% 94.4% 94.4% 88.9% 92.0% 94.0% 93.1% 93.8%
[…]
Univariate Design: WGAN DGP
Panel B: Vote Share (truth $\bm{=0.0992}$)
Bias 0.0013 0.0013 0.0013 0.0014 0.0016 -0.0031 -0.0045 -0.0042 -0.0033 -0.0025 0.0102 0.0023 0.0017 0.0000 0.0028 0.0055 0.0010 0.0015 -0.0001 0.0027
Variance 0.0007 0.0002 0.0001 0.0000 0.0000 0.0002 0.0000 0.0000 0.0000 0.0000 0.0009 0.0008 0.0010 0.0008 0.0008 0.0015 0.0012 0.0012 0.0008 0.0008
95% coverage 94.1% 93.5% 95.5% 93.9% 91.6% 95.7% 93.2% 92.5% 76.6% 72.7% 95.7% 96.1% 94.2% 96.6% 95.2% 93.1% 94.1% 92.9% 96.2% 95.2%
[…]
Bivariate Design: Polynomial DGP
Panel C: Turnout Probability (truth $\bm{=-0.0376}$)
Bias -0.0118 -0.0201 -0.0120 -0.0185 -0.0161 -0.0049 -0.0043 0.0013 -0.0011 -0.0009 0.0220 0.0134 0.0117 0.0078 0.0011 0.0169 0.0127 0.0106 0.0079 0.0012
Variance 0.0786 0.0127 0.0051 0.0010 0.0005 0.0189 0.0060 0.0034 0.0011 0.0007 0.0070 0.0063 0.0060 0.0061 0.0060 0.0084 0.0064 0.0060 0.0059 0.0058
95% coverage 93.4% 94.8% 95.3% 93.1% 93.3% 97.8% 97.3% 98.1% 98.3% 97.1% 95.3% 94.9% 95.2% 94.7% 93.6% 94.1% 94.9% 94.8% 94.0% 93.7%
Panel D: Housing Price (truth $\bm{=1.4878}$)
Bias -3.8353 -4.0482 -3.1878 -1.9335 -1.2293 0.8840 -0.2467 -0.5726 -0.2619 -0.1799 6.7977 3.9785 3.4101 1.9859 2.1574 1.2121 0.8221 0.8680 0.0998 0.6290
Variance 528.8758 85.8068 37.4150 9.2601 5.7001 315.3792 95.7504 59.4647 15.2623 8.8313 49.9737 37.6548 39.3580 34.7354 37.3014 87.1682 59.4337 57.0010 46.7899 48.0771
95% coverage 94.8% 94.7% 95.5% 95.7% 95.0% 97.2% 97.8% 97.4% 97.6% 97.0% 82.4% 92.4% 91.5% 94.1% 91.9% 93.8% 94.6% 95.8% 93.2% 93.6%
Panel E: Age (truth $\bm{=1.0880}$)
Bias -0.0178 -0.1874 -0.1966 -0.1698 -0.1408 -0.2898 0.0052 0.0626 0.0390 -0.0502 -0.2220 -0.2153 -0.2951 0.0170 -0.2594 0.2984 -0.0097 -0.1705 0.0631 -0.2111
Variance 110.6967 15.8490 7.9903 1.2539 0.6609 36.1169 9.6130 6.1630 2.1915 1.2345 9.2555 9.1571 8.4339 7.8535 7.5993 13.8381 11.5593 9.9945 8.5617 8.2755
95% coverage 93.3% 94.9% 94.2% 97.1% 95.9% 97.6% 97.6% 97.7% 97.4% 96.8% 95.7% 93.9% 94.5% 94.2% 94.7% 94.6% 93.5% 93.6% 93.7% 94.9%
[…]
Bivariate Design: WGAN DGP
Panel F: Turnout Probability (truth $\bm{=0.0356}$)
Bias 0.0180 0.0294 0.0257 0.0259 0.0252 -0.0061 -0.0063 -0.0022 -0.0019 -0.0020 0.0490 0.0309 0.0242 0.0211 0.0177 0.0426 0.0280 0.0217 0.0197 0.0169
Variance 0.2461 0.0240 0.0094 0.0015 0.0008 0.0305 0.0122 0.0075 0.0028 0.0015 0.0054 0.0046 0.0043 0.0039 0.0040 0.0069 0.0051 0.0045 0.0036 0.0037
95% coverage 94.8% 93.6% 93.8% 90.1% 85.5% 98.5% 97.9% 98.4% 97.9% 99.1% 90.7% 93.8% 93.8% 94.9% 93.9% 86.8% 90.6% 93.1% 94.4% 93.0%
Panel G: Housing Price (truth $\bm{=4.0414}$)
Bias 2.7574 4.8878 5.6543 4.5000 3.8748 -0.2310 -0.5278 -0.0903 -0.1227 -0.0275 8.7037 5.7927 4.9030 2.4293 2.1113 6.3921 4.7853 4.2545 2.2442 1.9833
Variance 3806.2308 232.4295 80.8610 10.4512 5.0567 423.0745 180.9454 109.4033 29.0582 14.4899 37.0680 34.4074 31.7393 27.7547 25.9484 72.5886 50.0642 44.7507 30.8009 27.3207
95% coverage 94.0% 93.4% 93.0% 88.0% 84.0% 98.1% 97.4% 97.2% 97.9% 98.7% 72.2% 81.5% 85.3% 91.9% 92.2% 80.6% 84.3% 85.5% 91.0% 93.4%
Panel H: Age (truth $\bm{=4.4519}$)
Bias -2.2493 -1.2546 -1.1629 -1.2299 -1.2656 -0.2949 0.0769 0.0839 -0.0428 -0.0237 -0.6848 -0.1871 -0.1209 -0.0749 -0.2052 -0.7206 -0.1786 -0.0985 -0.0659 -0.2074
Variance 2200.4077 58.0355 21.9401 2.8652 1.4544 68.6196 23.8051 14.7071 4.4742 2.4230 8.2230 7.2583 6.4681 6.4340 5.8234 13.4557 8.5387 7.0771 6.5556 5.9329
95% coverage 92.1% 94.4% 94.5% 91.7% 85.1% 98.1% 98.3% 98.8% 98.4% 98.3% 96.0% 95.0% 95.3% 94.3% 94.6% 93.8% 93.8% 94.9% 94.9% 95.0%
Table 1: Monte Carlo simulation results. For each method, we vary the sample size $n\in\{1000,5000,10000,50000,100000\}$ and report the average bias, variance, and $95\%$ coverage rate across 1,000 Monte Carlo replications. Each panel corresponds to a different outcome variable in Section 4.1, with the true treatment effect in parentheses either calculated exactly (from the polynomial fit) or simulated as described in Section 4.2 (for the WGAN DGPs).
Local linear regressions perform well in univariate settings, although they slightly undercover in the polynomial DGP (Panel A, Columns 1-5 of Table 1). [^9] This undercoverage is consistent with the simulation result ($90.6\%$ coverage) of Calonico et al. (2019), where they use the same polynomial DGP but with $n=1000$ for a total of $5000$ replications; see Table SA-1, Row 2 and Column 3 in their supplemental appendix. We have done our simulation using their STATA code and obtained similar results to those reported in Table 1. However, they are in general noisier in bivariate settings, where we follow Keele and Titiunik (2015) to first collapse the bivariate score using rescaled Euclidean (chordal) distance and then run rdrobust, especially in small samples. Although the variance of local linear regressions shrinks as the sample size gets large, their bias does not shrink fast enough relative to the diminishing variance in those bivariate WGAN DGPs (Panels F-H, Columns 1-5 of Table 1), and therefore the coverage rate drops as the sample size increases. This problem arises possibly due to the zero-density issue of the collapsed multivariate score at the new univariate cutoff, as we point out in Section 3.1.1 and empirically illustrate in the last two histograms of Figure 1, cautioning against univariate transformations of multivariate scores.
The minimax-optimal estimator of Imbens and Wager (2019) has a relatively low bias but tends to overcover, which is expected since it is designed for minimizing the worst-case MSE and the construction of confidence intervals aims at achieving uniform coverage over a class of CEFs explained in Section 3.2. Nevertheless, its performance depends crucially on accurate estimates of the second derivative bound $\mathcal{B}$: for the two artificial datasets built using the Lee (2008) data (Panels A-B of Table 1), the minimax-optimal estimator undercovers severely, especially for the polynomial DGP in Panel A. There, we follow the recommendation of Imbens and Wager (2019) to estimate the CEF globally using a second-order polynomial and multiply the maximal estimated curvature by $4$. But this turns out to yield a too small estimate of $\mathcal{B}$; Table B.1 in Appendix B.2 shows how the simulation results change with the constant we multiply, and it is only after multiplying the maximal estimated curvature by $30$ do we get an estimate that is close to the true (least upper) bound $\mathcal{B}$. Of course, one could argue that the global second-order polynomial is a bad model to start with; in the bivariate design, we follow their working paper (Imbens and Wager, 2018, p.21) and estimate $\mathcal{B}$ using the $95$-th percentile (over all sample points) of the operator norm curvature via a cross-validated ridge regression with interacted $7$-th order natural splines, which yields good coverage results (Panels C-H, Columns 6-10 of Table B.1). However, without knowledge about the true DGP, choosing a “good” model is not a straightforward problem, and estimating second derivatives is more difficult when the dimension of the score increases. In addition, the current implementation of optrdd accepts up to two scores only, limiting its application more broadly.
Honest regression forests and local linear forests have comparable performance to local linear regressions in the univariate design (Panels A-B, Columns 11-20 of Table 1), though their variances are slightly higher. On the other hand, regression forests are in general more biased than local linear forests, especially in small samples, which is a known boundary problem for regression forests. In addition, the variance of regression forests tends to be smaller than that of local linear forests, exacerbating the undercoverage problem. As we discussed in Section 3.3.2, local linear forests are better at capturing smooth signals via a linear correction, and its advantage over regression forests is especially pronounced when the underlying DGP exhibits strong linear trends near the boundary point at which we estimate the treatment effect. In Appendix B.5, we examine the robustness of the results in Table 1 to the number of training epochs, which is a tuning parameter of WGAN, and increasing the number of training epochs leads to stronger linear trends near the boundary point (see Figure B.7); in these cases, the performance of regression forests deteriorates, whereas local linear regressions have robust performance. With that said, these two forest-based methods are valid alternatives to existing methods for multivariate RD designs that can flexibly model any given number of scores, and empirical researchers may leverage and tailor them accordingly to fit different empirical settings.
5 Conclusion
This paper studies RD designs with multiple scores. We first highlight the critical but often overlooked zero-density issue with running local linear regressions using a collapsed univariate score, and then explore the use of nonparametric estimators based on forests that directly model multivariate scores. Our simulation results suggest that the zero-density issue, while hard to detect from an empirical dataset, is likely present and would impair the performance of local linear regressions. Regression forests and local linear forests, on the other hand, generally require larger sample sizes to achieve good performance at boundary points, though local linear forests are better at capturing linear trends. Although this paper does not offer a simple, universal solution to multivariate RD problems, it adds two forest-based estimators to empirical researchers’ toolkit. Overall, we recommend using the two forest-based methods as the primary approach when the sample size is large and the number of score is moderate because of their flexibility and theoretical validity. In cases where the sample size is small, local linear regressions with a collapsed univariate score can serve as a starting point, but researchers should proceed with caution and perhaps compare the results against other nonparametric alternatives introduced in this paper as a way of robustness checks.
SUPPLEMENTARY MATERIAL
Appendix A Proof
Proof of Proposition 1. Define $Z=(X_{2}-c_{2})^{2}$. Consider the mapping $(X_{1},Z)\mapsto(E,V)$, which is one-to-one. We can find the joint density of $(X_{1},Z)$ by
$\mathbb{P}(X_{1}\leq x_{1},Z\leq z)$ $=\mathbb{P}\left(X_{1}\leq x_{1},c_{2}-\sqrt{z}\leq X_{2}\leq c_{2}+\sqrt{z}\right)$
$=F_{(X_{1},X_{2})}\left(x_{1},c_{2}+\sqrt{z}\right)-F_{(X_{1},X_{2})}\left(x_{1},c_{2}-\sqrt{z}\right)$
$\implies f_{(X_{1},Z)}(x_{1},z)$ $=\frac{\partial^{2}{\mathbb{P}(X_{1}\leq x_{1},Z\leq z)}}{\partial x_{1}\partial z}$
$=\frac{1}{2\sqrt{z}}\left[f_{(X_{1},X_{2})}\left(x_{1},c_{2}+\sqrt{z}\right)+f_{(X_{1},X_{2})}\left(x_{1},c_{2}-\sqrt{z}\right)\right].$
Since $X_{1}=V+c_{1}$ and $Z=E^{2}-V^{2}$, we have the Jacobian
$J_{(X_{1},Z)}(e,v)=\begin{bmatrix}0&1\\ 2e&-2v\end{bmatrix}$
and $\lvert J_{(X_{1},Z)}(e,v)\rvert=2e$. By the bivariate transformation theorem,
$f_{(E,V)}(e,v)$ $=f_{(X,Z)}(x,z)\lvert J_{(X_{1},Z)}(e,v)\rvert$
$=\frac{e}{\sqrt{e^{2}-v^{2}}}\left[f_{(X_{1},X_{2})}\left(v+c_{1},c_{2}+\sqrt{e^{2}-v^{2}}\right)+f_{(X_{1},X_{2})}\left(v+c_{1},c_{2}-\sqrt{e^{2}-v^{2}}\right)\right].$
This implies that the marginal distribution of $E$ is
[…]
and the claim follows. $\square$
A.1 Examples
Example A.1. Let $(X_{1},X_{2})$ be uniformly distributed over $[-1,1]^{2}$. Then $f_{(X_{1},X_{2})}(x_{1},x_{2})=\frac{1}{4}$ over the support $\Omega(X_{1},X_{2})=[-1,1]^{2}$. First consider the boundary point at $(0,0)$. The bivariate transform $(E,V)$ in this case has joint support
$\Omega(E,V)$ =$=\left\{(e,v)\in\mathbb{R}^{2}:e\in[0,\sqrt{2}]\right.,$
$\left.\qquad\qquad v\in[-e,e]\text{ if }e\in[0,1]\text{ and }v\in\left[\sqrt{e^{2}-1},1\right]\cup\left[-1,-\sqrt{e^{2}-1}\right]\text{ if }e\in\left(1,\sqrt{2}\right]\right\}.$
Then
[…]
Following Proposition 1,
$f_{E}(e)=$ $\pi e/2,$ if $e\in[0,1]$
$f_{E}(e)=$ $e\left[\sin^{-1}(1/e)-\sin^{-1}(\sqrt{e^{2}-1}/e)\right],$ if $e\in(1,\sqrt{2}]$. (18)
Hence $f_{E}(0)=0$. Note that this is not specific to the chosen boundary point, $(0,0)$. In fact, the issue remains for any boundary point $(c_{1},c_{2})\in[-1,1]^{2}$ with respect to which we calculate the Euclidean distance. To see this, suppose without loss of generality that $c_{1},c_{2}>0$; other cases follow similarly. Then
$E=\sqrt{(X_{1}-c_{1})^{2}+(X_{2}-c_{2})^{2}},\quad V=X_{1}-c_{1},$
where
$|V|\leq E\leq\sqrt{V^{2}+(1+c_{2})^{2}},\quad E\in\left[0,\sqrt{(1+c_{1})^{2}+(1+c_{2})^{2}}\right].$
Then the joint support of $(E,V)$ when $E$ takes value in $[0,1-c_{1}]$ is
$\Omega\left(E,V\,|\,E\in[0,1-c_{1}]\right)=\left\{(e,v)\in\mathbb{R}^{2}:e\in[0,1-c_{1}],v\in[-e,e]\right\}.$
The case where $E\in\left(1-c_{1},\sqrt{(1+c_{1})^{2}+(1+c_{2})^{2}}\right]$ is irrelevant to deriving $f_{E}(0)$. Then Equation (18) still holds for $e\in[0,1-c_{1}]$ and therefore $f_{E}(0)=0$.
Example A.2. Let $(X_{1},X_{2})$ be independently normally distributed with mean 00 and variance $\sigma^{2}$. Then $f_{(X_{1},X_{2})}(x_{1},x_{2})=\frac{1}{2\pi\sigma^{2}}\exp\{-x_{1}^{2}/2\sigma^{2}\}\exp\{-x_{2}^{2}/2\sigma^{2}\}$ over the support $\Omega(X_{1},X_{2})=\mathbb{R}^{2}$. First consider the boundary point at $(0,0)$. The bivariate transform $(E,V)$ in this case has joint support
$\Omega(E,V)$ =$=\left\{(e,v)\in\mathbb{R}^{2}:e\geq 0,v\in[-e,e]\right\}.$ (19)
Then
[…]
$=\frac{e}{\sigma^{2}}\exp\{-e^{2}/2\sigma^{2}\}$
Hence $f_{E}(0)=0$. Note that this is again not specific to the choice of the boundary point, as for any $(c_{1},c_{2})\in\mathbb{R}^{2}$, Equation (19) still holds.
Appendix B More on Simulations
This section provides additional information on the Monte Carlo simulations in Section 4.
B.1 Details on the Polynomial DGPs
B.1.1 Univariate Design: Lee (2008) Data
Following Imbens and Kalyanaraman (2012), Calonico et al. (2014) and Calonico et al. (2019), we build the polynomial DGP using the Lee (2008) data as follows:
$Y_{i}=\mu(X_{i})+\epsilon_{i},\quad X_{i}\sim(2{Beta}(2,4)-1),\quad\epsilon\sim N(0,\sigma^{2})$
where ${Beta}$ denotes the Beta distribution, $\mu(\cdot)$ is constructed by fitting a $5$-th order global polynomial with different coefficients for the treated and untreated groups taking the form:
$\mu(x)=\begin{cases}0.48+1.27x+7.18x^{2}+20.21x^{3}+21.54x^{4}+7.33x^{5}&\text{if $x<0$}\\ 0.52+0.84x-3.00x^{2}+7.99x^{3}-9.01x^{4}+3.56x^{5}&\text{if $x\geq 0$}\end{cases}$ (20)
and $\sigma$ is estimated from the residuals of the regression and is equal to $0.1295$. Under this DGP, the true treatment effect at the 00 cutoff is $0.04$.
B.1.2 Bivariate Design: Keele and Titiunik (2015) Data
For the continuous outcome variables, housing price and age, the models are of the form
${Y_{i}}_{j}=\mu_{j}(X_{i})+\epsilon_{ij},\text{ where }j\in\{\text{price},\text{age}\}\text{ and }\epsilon_{ij}\sim N(0,\sigma_{j}^{2}).$
We fit third-order interacted polynomials for $\mu_{j}(X_{i})$. Below are the model specifications, where $\tilde{x}_{i1},\tilde{x}_{i2}$ are demeaned versions of $x_{i1},x_{i2}$ (i.e., latitude, longitude) respectively:
$\mu_{price}(x_{i})=\begin{cases}13544.3+150.2{x}_{i1}-11847.6\tilde{x}_{i1}^{2}-29821.7\tilde{x}_{i1}^{3}\\ \quad+259.3{x}_{i2}-15479.8\tilde{x}_{i2}^{2}-207469.6\tilde{x}_{i2}^{3}+24446.9\tilde{x}_{i1}\tilde{x}_{i2}&\text{if $i$ is treated,}\\ 230407.6-9713.5{x}_{i1}+537644.5\tilde{x}_{i1}^{2}-8356369.6\tilde{x}_{i1}^{3}\\ \quad-2164.6{x}_{i2}-9998.4\tilde{x}_{i2}^{2}+748352.9\tilde{x}_{i2}^{3}+55631.1\tilde{x}_{i1}\tilde{x}_{i2}&\text{if $i$ is control,}\end{cases}$
and $\sigma_{price}=32.6334$.
$\mu_{age}(x_{i})=\begin{cases}-477.8-70.0{x}_{i1}-111.9\tilde{x}_{i1}^{2}-54704.1\tilde{x}_{i1}^{3}\\ \quad-44.9{x}_{i2}+4564.6\tilde{x}_{i2}^{2}+107699.6\tilde{x}_{i2}^{3}-5863.3\tilde{x}_{i1}\tilde{x}_{i2}&\text{if $i$ is treated,}\\ 77307.5-1026.2{x}_{i1}+60847.1\tilde{x}_{i1}^{2}-749930.1\tilde{x}_{i1}^{3}\\ \quad+481.1{x}_{i2}-13224.5\tilde{x}_{i2}^{2}+185693.5\tilde{x}_{i2}^{3}-16725.5\tilde{x}_{i1}\tilde{x}_{i2}&\text{if $i$ is control,}\end{cases}$
and $\sigma_{age}=15.9496$. For the discrete outcome, election turnout, we consider a logit conditional expectation function $P(Y_{i}=1|X_{i})=\frac{e^{poly(X_{i})}}{1+e^{poly(X_{i})}}$, where $poly(X_{i})$ is a third-order interacted polynomial as before. Below is the specification for $poly(X_{i})$:
$poly(x_{i})=\begin{cases}-200.4+3.7{x}_{i1}+224.4\tilde{x}_{i1}^{2}+1028.1\tilde{x}_{i1}^{3}\\ \quad-0.7{x}_{i2}+57.1\tilde{x}_{i2}^{2}+3778.1\tilde{x}_{i2}^{3}-127.5\tilde{x}_{i1}\tilde{x}_{i2}&\text{if $i$ is treated,}\\ 10399.6-286.5{x}_{i1}+10516.0\tilde{x}_{i1}^{2}-114884.8\tilde{x}_{i1}^{3}\\ \quad-15.4{x}_{i2}-423.8\tilde{x}_{i2}^{2}+12700.1\tilde{x}_{i2}^{3}+228.7\tilde{x}_{i1}\tilde{x}_{i2}&\text{if $i$ is control.}\end{cases}$
Figure B.1 visualizes the polynomial and WGAN DGPs for age and housing price.
Figure B.1: Visualization of the sample data in Keele and Titiunik (2015) (left in each row) as well as simulated polynomial and WGAN DGPs (middle and right in each row, respectively). The outcome variable for the first (resp. second) row is housing price (resp. age). Coordinates below the black curve (treatment boundary) are treated, and the green circle shows the boundary point at which we estimate the treatment effect. The plot for actual housing prices uses a different dataset, which has fewer observations than the other two outcome cases. To visualize the polynomial DGPs, 2,500 coordinates are sampled with replacement from the original sample plus a noise term $N(0,0.01^{2})$. We fit $3$-rd order interacted polynomials for housing price and age, and the plots show the model predictions at each pair of coordinates sampled.
B.2 Estimating the Second Derivative Bound
In this section, we consider estimating the second derivative bound $\mathcal{B}$ for the univariate polynomial DGP based on the real data from Lee (2008) by fitting a global quadratic function and multiplying the maximal estimated curvature by different constants. The results are reported in Table B.1, along with the results using the infeasible least upper bound $\mathcal{B}$ calculated from the true polynomial DGP in (20). Imbens and Wager (2019) recommend multiplying by $2$ or $4$, but in this example these choices do not yield reasonable estimates for the second derivative bound and thus lead to biased estimates and poor coverage rates, as in Panels A-B, Columns 6-10 of Table 1. Only after multiplying the maximal estimated curvature by $30$ do we get a $\hat{\mathcal{B}}$ close to the true $\mathcal{B}$.
True $\mathcal{B}=14.36$$\hat{\mathcal{B}}=2\times$max.curvature ^$\hat{\mathcal{B}}=4\times$max.curvature
Sample size $n=$ $1,000$ $5,000$ $10,000$ $1,000$ $5,000$ $10,000$ $1,000$ $5,000$ $10,000$
Bias 0.0538 0.0471 0.0453 0.0432 0.0345 0.0297 0.0364 0.0252 0.0211
Variance 0.0019 0.0006 0.0003 0.0006 0.0002 0.0001 0.0009 0.0002 0.0002
95% coverage 96.7% 96.7% 97.0% 67.6% 36.7% 24.7% 80.3% 69.7% 64.0%
Average $\hat{\mathcal{B}}$ 0.8372 0.8363 0.8369 1.6555 1.6736 1.6791
^$\hat{\mathcal{B}}=6\times$max.curvature ^$\hat{\mathcal{B}}=10\times$max.curvature ^$\hat{\mathcal{B}}=30\times$max.curvature
Sample size $n=$ $1,000$ $5,000$ $10,000$ $1,000$ $5,000$ $10,000$ $1,000$ $5,000$ $10,000$
Bias 0.0306 0.0203 0.0166 0.0241 0.0145 0.0120 0.0115 0.0067 0.0056
Variance 0.0011 0.0003 0.0002 0.0013 0.0004 0.0002 0.0020 0.0005 0.0003
95% coverage 87.6% 85.0% 80.3% 93.0% 90.1% 89.8% 95.9% 96.5% 96.4%
Average $\hat{\mathcal{B}}$ 2.5116 2.5161 2.5217 4.1695 4.1910 4.1817 12.6209 12.5831 12.5993
Table B.1: Simulation results based on different estimates of the second derivative bound. The true least upper bound on the second derivative of the DGP (20) is $\mathcal{B}=14.36$. The table reports the average bias, variance, and $95\%$ coverage rate over the 1,000 Monte Carlo replications using the true bound as well as estimated bounds by fitting a global quadratic and then multiplying the maximal curvature of the fitted model by different constants $\in\{2,4,6,10,30\}$. Also reported is the Monte Carlo average of estimated bounds.
B.3 Details on the WGAN DGPs
To construct artificial datasets that closely resemble the observed samples in Keele and Titiunik (2015) near the treatment boundary, we employ the WGAN approach in Athey et al. (2021) and constrain the input data to a window defined by the start and end points of the treatment boundary, i.e., using the boundary as a rough diagonal. This window and the sample data points involved are represented by the plots on the left in Figures 3 and B.1. The resulting input data consists of 2,628 treated and 2,969 control observations for turnout and age, and 117 treated and 206 control observations for housing price.
We use the default options available from the wgan package of Athey et al. (2021) to train the WGAN for the results reported in Table 1.101010These default specifications can be found at ds-wgan.readthedocs.io/en/latest/api.html#specifications. We note that the current implementation of WGAN is only available in Python, whereas implementations of the minimax-optimal estimator and the two forests estimators are in R, and therefore we cannot, for each Monte Carlo replication, generate a new sample from WGAN in Python, keep the generated sample in memory, and then take it to R and fit the estimators: for one iteration of $1000$ Monte Carlo replications and each with a sample size of $n$, such procedure requires $n\times 1000$ observations, which quickly runs out of computer memory as we increase $n$. Instead, for each dataset that we fit a WGAN to, we draw a WGAN-generated sample of $40$ million observations, and use this large artificial sample to run the Monte Carlo simulations, treating it as a “superpopulation” from which we draw random subsamples of size $n$ used in each Monte Carlo replication. [^11] Each WGAN sample is corrected based on the geographic coordinates of the treatment boundary to ensure that for each hypothetical household, the generated coordinates match the generated treatment status. This correction is minor, affecting fewer than 2% of the observations in the WGAN sample. Figures B.2B.6 display graphical comparisons between the actual turnout, housing price and age samples and their WGAN-generated counterparts.
Figure B.2: Marginal histograms for the actual turnout and age samples (real) and their WGAN-generated counterparts (fake). The two observed samples share the same geographic coordinates, so the latitude and longitude comparisons are produced once (two plots on the left).
Figure B.3: Marginal histograms for the actual housing price sample (real) and its WGAN-generated counterpart (fake). The geographic coordinates in the observed price sample are a subset of those in the turnout/age sample, so the latitude and longitude comparisons are regenerated, while the WGAN coordinates are generated once and fixed across the three outcomes.
Figure B.4: Between-variable correlations for the observed turnout sample (real) and its WGAN-generated counterpart (fake).
Figure B.5: Between-variable correlations for the observed housing price sample (real) and its WGAN-generated counterpart (fake).
Figure B.6: Between-variable correlations for the observed age sample (real) and its WGAN-generated counterpart (fake).
B.4 Wasserstein Distances between Real and Artificial Data
Athey et al. (2021) combine generative adversarial networks (GANs) with the Wasserstein distance objective to create a hypothetical dataset that closely mimics the observed sample, enabling effective Monte Carlo simulations. In this subsection, we consider the two-dimensional example in Keele and Titiunik (2015) and show that the datasets constructed by WGAN are indeed more similar to the actual samples in terms of Wasserstein distance, as compared to the datasets generated by the polynomial DGPs in Section B.1.2. While we use the default options when training the WGAN DGPs used in the main text (training the neural network for 1,000 epochs, for example), we report the Wasserstein distances for 2,000, 3,000, and 4,000 epochs as well. As elaborated in Section B.5.3, these specifications serve as robustness checks for the estimators proposed in the main text.
Following Athey et al. (2021), we compute the exact Wasserstein distances via linear programming, averaging over 10 random samples from the WGAN dataset, each of equal size to the original dataset. Table B.2 organizes the Wasserstein distances for the outcome variables considered in Section B.3. We note that the Wasserstein distance between distributions $\mathbb{P}$ and $\mathbb{P^{\prime}}$, defined as
$W(\mathbb{P},\mathbb{P^{\prime}}):=\inf_{\gamma\in\Pi(\mathbb{P},\mathbb{P^{\prime}})}\mathbb{E}_{(R,S)\stackrel{{\scriptstyle d}}{{\sim}}\gamma}\left[||R-S||\right]$
with $\Pi(\mathbb{P},\mathbb{P^{\prime}})$ denoting the set of joint distributions with marginals $\mathbb{P}$ and $\mathbb{P^{\prime}}$ such that $R\stackrel{{\scriptstyle d}}{{\sim}}\mathbb{P}$ and $S\stackrel{{\scriptstyle d}}{{\sim}}\mathbb{P^{\prime}}$, is not normalized with respect to the scale of random variables $R$ and $S$. Thus, its magnitude can depend on the scale of the data being compared. This explains why housing price has the largest Wasserstein distances and turnout has the smallest. Comparing across rows in Table B.2, all the WGAN DGPs considerably outperform the polynomial ones in terms of Wasserstein distance. Within the WGAN entries, the distances for turnout and age tend to shrink with increasing training epochs. However, for housing price, the small sample size (117 treated and 206 control observations) may explain the instability in training the networks and calculating the distances.
Polynomial WGAN
1,000 Epochs 2,000 Epochs 3,000 Epochs 4,000 Epochs
Turnout 142.8712 0.1603 0.0147 0.0242 0.0047
Housing Price 330.3712 3.6796 6.5035 6.1220 7.0828
Age 147.0418 1.9824 1.1613 1.3046 1.0996
Table B.2: Wasserstein distances between generated and observed data.
B.5 Robustness of the WGAN simulations
For the bivariate designs in Table 1, results on the housing price outcome appear to be most sensitive to changes in sample size and estimator choice. To complement Section 4.3, we consider three robustness checks: robustness to subsamples, model architectures, and the number of training epochs. The first two checks are also considered by Athey et al. (2021) to assess the ranking of a set of estimators for the average effect for the treated. We omit the third robustness check in Athey et al. (2021)—robustness to the size of training data—since the sample size of the housing price data (117 treated and 206 control observations) is too small for most of the fractions considered in Athey et al. (2021, Table 11).
B.5.1 Robustness to Subsamples
Following Athey et al. (2021, Section 5.1), we run the WGAN algorithm on 10 subsamples randomly drawn without replacement from the original data; each subsample is 80% of the size of the original data. [^12] More specifically, the WGAN coordinates ($X_{i}$) and treatment statuses ($D_{i}$) are generated from 80% of the turnout/age data, and the corresponding housing prices ($Y_{i}|(X_{i},D_{i})$) are generated from 80% of the housing price data. This is because the coordinates in the housing price data are a subset of those in the turnout/age data, which has more observations. For each subsample, we train a WGAN superpopulation of size 5 million and run 1,000 Monte Carlo simulations with sample size 10,000 for each of the four estimators to estimate the treatment effect at the boundary point. [^13] We consider a smaller superpopulation size because we need to train a WGAN for each of the subsamples drawn, and repeatedly generating a superpopulation of size 40 million as in Section 4 turned out to be too computationally intensive. Table B.3 shows the bias, variance and coverage averaged across the 10 hypothetical populations trained from the 10 random subsamples of the original data. Standard deviations are enclosed in parentheses. The averaged metrics from using the subsamples do not differ considerably from those from using the full sample, see Table 1, Panel G under $n=$ 10,000.
Local Linear Regression Minimax-Optimal Honest Regression Forest Honest Local Linear Forest
Bias 5.2632 (1.4239) 0.0176 (0.5703) 4.7418 (1.3558) 3.5876 (0.9605)
Variance 82.4645 (25.0978) 106.13 (42.6781) 28.7466 (6.2132) 42.3938 (12.0164)
95% coverage 94.22% (0.71%) 98.07% (0.59%) 84.88% (5.71%) 88.43% (3.05%)
Table B.3: Robustness results for housing price: average and standard deviations of metrics over 10 subsamples drawn from the original sample. Each random subsample for generating the WGAN superpopulation is 80% of the original sample size. The sample size is 10,000 for each Monte Carlo trial.
B.5.2 Robustness to Model Architectures
The default architecture in the WGAN algorithm for the generator and critic neural networks consists of three hidden layers with 128 units each. To examine whether our qualitative comparisons in Section 4.3 are sensitive to the architecture specification, we adopt two alternative model architectures from Athey et al. (2021, Section 5.2). Both alternatives have a three-layer generator and a three-layer critic. In the first alternative, the generator (critic) has layer dimensions of $[64,128,256]$ ($[256,128,64]$). In the second alternative, the generator (critic) has layer dimensions of $[128,256,64]$ ($[64,256,128]$). Both WGAN superpopulations are of size 5 million, and the sample size is 10,000 for each Monte Carlo replication. Results from varying the model architecture are organized in Table B.4.
Local Linear Regression Minimax-Optimal Honest Regression Forest Honest Local Linear Forest
Main (generator: $[128,128,128]$, critic: $[128,128,128]$)
Bias 5.6543 $-0.0903$ 4.9030 4.2545
Variance 80.8610 109.4033 31.7393 44.7507
95% coverage 93.0% 97.2% 85.3% 85.5%
Alternative 1 (generator: $[64,128,256]$, critic: $[256,128,64]$)
Bias 4.6898 0.1933 2.9349 2.3774
Variance 60.4756 88.1598 23.5457 31.2027
95% coverage 95.0% 97.8% 90.6% 91.4%
Alternative 2 (generator: $[128,256,64]$, critic: $[64,256,128]$)
Bias 6.1217 0.3147 6.6993 5.5907
Variance 97.7364 155.9217 31.0610 48.2774
95% coverage 93.3% 97.3% 78.4% 81.8%
Table B.4: Robustness results for housing price, with the first (second) alternative architecture having a generator hidden layer with dimensions $[64,128,256]$ ($[128,256,64]$) and a critic hidden layer with dimensions $[256,128,64]$ ($[64,256,128]$). The sample size 10,000.
B.5.3 Robustness to the Number of Training Epochs
In this subsection, we vary the number of training epochs specified in the WGAN algorithm for the housing price outcome as a robustness check. We consider running the algorithm for 2,000, 3,000, and 4,000 epochs, noting that 1,000 is the default option used for constructing the datasets in the main text. Each Monte Carlo trial uses a sample size of 10,000, which is moderate-to-large for RD studies. Figure B.7 shows three-dimensional surface plots that visualize the simulated CEF of housing price from the polynomial and the four WGAN DGPs, with the original observations scattered around the surfaces (i.e., CEFs). While the WGAN surfaces generally exhibit less curvature than the polynomial ones, we point out that the curvature of the treated WGAN surface (surrounded by treated dots in red) tends to increase with the number of training epochs. This instability may be attributed to the small sample size of the original data, which includes only 117 treated observations. Nevertheless, such variability in curvature could highlight certain strengths and weaknesses of the four estimators discussed in the main text. Table B.5 presents the simulation results using the three alternative DGPs. Similar to the 1,000-epoch DGP in the main text, for $n=$ 10,000, local linear regressions perform well, and the minimax-optimal estimator overcovers. [^14] Although not shown in Table B.5, we note that consistent with Panel G in Table 1, the coverage rates for local linear regressions again deteriorate as we further increase the sample size. This deterioration is possibly due to the zero-density issue of the collapsed multivariate score. Honest regression forests, however, gradually fail to capture linear signals in the treated surface near the boundary point as the surface becomes more tilted. Local linear forests alleviate this problem through the inclusion of a linear correction, as shown by their robust performance across DGPs. Nevertheless, forest-based estimators generally require larger sample sizes to achieve good coverage rates.
Figure B.7: Surface plots for simulated housing price using the polynomial DGP detailed in Section B.1.2 and the WGAN DGP introduced in Section 4.2 with 1,000, 2,000, 3,000 and 4,000 training epochs. Each point on the surfaces represents the average of 2,000 simulated housing prices at the corresponding geographic location. The outcome at the focal point is evaluated both as treated (solid red dot) and not treated (solid black dot), also averaged over 2,000 simulations. The original sample is displayed as a scatter plot around the surfaces, with red (treated) and black (control) semi-transparent dots.
Local Linear Regression Minimax-Optimal Honest Regression Forest Honest Local Linear Forest
2,000 Epochs
Bias 2.2475 $-0.8902$ 6.8735 3.4928
Variance 190.9314 142.4837 24.5066 71.0147
95% coverage 94.3% 99.4% 73.8% 85.0%
3,000 Epochs
Bias $-3.2831$ $-1.8100$ 10.7442 4.4552
Variance 304.8468 122.9466 21.3419 82.1174
95% coverage 93.9% 99.3% 40.5% 79.9%
4,000 Epochs
Bias $-3.4504$ $-1.2560$ 11.3784 3.1766
Variance 258.7163 151.0802 24.7439 103.9739
95% coverage 93.7% 98.4% 39.0% 81.4%
Table B.5: Robustness results for housing price, with {2,000, 3,000, 4000} training epochs in the WGAN specification. The sample size 10,000.
B.6 Hyperparameters for Regression and Local Linear Forests
For all forest-based estimators, we set the number of trees (parameter B of the regression_forest() function in the grf package) to 5,000. More trees generally reduce overfitting but at a higher computational cost. At each sufficiently large tree node, we randomly pick one feature to make the split ($\texttt{mtry}=1$). Note that in univariate designs, mtry is always 1. Small values of mtry can reduce the between-tree correlations, which would lower the overall variance. For regression forests, our choice of the subsample fraction is based on the condition in Athey et al. (2019, Theorem 5) that the subsample sizes $s_{1}$ and $s_{0}$, for the treated and control samples respectively, should satisfy $s_{1}\asymp n_{1}^{\beta}\,\,\,\text{and}\,\,\,s_{0}\asymp n_{0}^{\beta}$ for some $\beta$ lower bounded by $\beta_{min}^{RF}:=\left(1+\frac{\texttt{mtry}\times\log(1-\alpha)}{d\log(\alpha)}\right)^{-1}$, where $\alpha$ is the maximum imbalance of a split that defaults to $0.05$. Given this condition, we set the subsample fraction to $ceil[n_{j}^{\beta_{min}^{RF}}]/n_{j}$ times a constant $c\in[0.05,0.5]$ for $j\in\{0,1\}$. We leave the other tunable parameters of regression_forest() at their default values for simplicity and generalizability. [^15] See the manual for grf: regression_forest() and ll_regression_forest() for details.
Figures B.8 and B.9 show, for each DGP we consider with sample size $n=$ 5,000, how the squared bias, variance, MSE and the $95\%$ coverage rate across the 1000 Monte Carlo replications change with the scaling constant $c$ we use. As $c$ gets smaller, the bias problem and undercoverage become more pronounced, whereas the variance decreases. We recommend setting $c$ in the range of $[0.4,0.5]$ for small to moderate datasets; smaller values of $c$ can be considered with larger sample sizes. In this paper, the simulations for all forest-based estimators are conducted with the setting $c=0.4$.
Figure B.8: Comparisons of different choices of the scaling constant $c$ using the datasets simulated by the polynomial DGPs. The horizontal axis corresponds to $c\in\{0.2,0.3,0.4,0.5\}$ as well as the “Default” option, in which case the subsample fraction defaults to $0.5$ without depending on $c$, $\beta$, or $n$. The left vertical axis corresponds to values of the Monte Carlo bias squared (red), variance (blue), and MSE (green); the right vertical axis corresponds to values of the empirical $95\%$ coverage rate (black).
Figure B.9: Comparisons of different choices of the scaling constant $c$ using the datasets simulated by the WGAN DGPs. The horizontal axis corresponds to $c\in\{0.2,0.3,0.4,0.5\}$ as well as the “Default” option, in which case the subsample fraction defaults to $0.5$ without depending on $c$, $\beta$, or $n$. The left vertical axis corresponds to values of the Monte Carlo bias squared (red), variance (blue), and MSE (green); the right vertical axis corresponds to values of the empirical $95\%$ coverage rate (black).
For local linear forests, the lower bound for $\beta$ is instead given by Friedberg et al. (2020, Theorem 5): $\beta_{min}^{LLF}:=1-\left(1+\frac{d}{1.3\times\texttt{mtry}}\cdot\frac{\log(\alpha)}{\log(1-\alpha)}\right)^{-1}$. Following Friedberg et al. (2020), we make residual splits (enable.ll.split = TRUE) to minimize the corresponding prediction errors on ridge regression residuals, as opposed to standard splits that minimize prediction errors from employing leaf-wide averages. We standardize the ridge penalty by the covariance matrix (ll.split.weight.penalty = TRUE). Other tunable parameters of ll_regression_forest() are set to their default values, except the ridge penalty term $\lambda$ is automatically tuned by cross-validation.
Although we do not tune the parameters in regression_forest() and ll_regression_forest() for ease of computation and to demonstrate the general applicability of the default specifications, we note that tuning may improve performance in terms of MSE but worsen coverage rates. This occurs because tuning targets out-of-bag MSE and trades off between bias and variance, which can lead to more biased results but smaller variances, thereby affecting the coverage rates.